OpenStudy (anonymous):

need to find the solution so i can graph it y=-5x-2 y+2=-5x

OpenStudy (amistre64):

what does y equal?

OpenStudy (anonymous):

-5x-2 i know to put that in the second equation so that it would be (-5x-2)+2=-5x that is where i get stuck

OpenStudy (amistre64):

well, you got that part good (-5x - 2) + 2 = -5x -5x - 2 + 2 = -5x ; well, we should know that -2+2 = 0 leaving us with -5x = -5x ; for what values of x does -5x equal -5x?

OpenStudy (anonymous):

infinite solutions right?

OpenStudy (amistre64):

correct, in fact both equations are the same equation written in different forms

OpenStudy (anonymous):

ok thank you!

OpenStudy (amistre64):

yep, so you do know how to graph the line for: y=-5x-2 right? :)

OpenStudy (anonymous):

yep

OpenStudy (amistre64):

good ;)